if f(x)=(x-1)/(x+1),then f^-1 (1/t)=?
Answer by stanbon(75887)

if f(x)=(x-1)/(x+1),then f^-1 (1/t)=?
--------------
1st:: Interchange x and y to get:
x = (y-1)/(y+1)
2nd: Solve for "y"::
xy+x = y-1
xy-y = -(x+1)
y(x-1) = -(x+1)
y = -(x+1)/(x-1)
-----
Then::
f^-1(1/t) = -(1/t + 1)/(1/t -1)
----
f^-1(1/t) = -(1+t)/t / (1-t)/t
----
f^-1(1/t) = -(1+t)/(1-t)
----------------------------
Cheers,
Stan H.
-------------